OpenStudy (anonymous):

45% OF THE 300 SHOPPERS IN A SUPERMARKET BOUGHT MILK. HOW MANY SHOPPERS IN THE SUPERMARKERT MILK?

OpenStudy (anonymous):

TURN THE PERCENTAGE INTO A DECIMAL AND MULTIPLY BY THE NUMBER OF SHOPPERS. FOR EXAMPLE IF 50% BUY MILK THAT MEANS .50* 300 = 150 BOUGHT MILK

OpenStudy (anonymous):

THE SALE PRICE OF $24 WAS 75% OF THE REGULAR PRICE. THE SALE PRICE WAS HOW MANY DOLLARS LESS THEN THE REGULAR PRICE?

OpenStudy (anonymous):

DO THE SAME AS THE FIRST BUT THEN YOU NEED TO SUBTRACT THE SALE PRICE FROM THE TOTAL

OpenStudy (anonymous):

HOW? I HATE THESE PROBLEMS CAN I HAVE A EXAMPLE?

OpenStudy (anonymous):

45/100 = 135/300 135 of the shoppers purchased milk.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome. Proportions work best with these types of problems.

OpenStudy (anonymous):

ok! hey can u help with this one? The sale price of $24 was 75%. the sale price was how many dollars less then the regular price?

OpenStudy (anonymous):

dont just give the answer, try to walk them through it.

OpenStudy (anonymous):

so you want to find the new sale price first. to do that turn the percentage into a decimal. to find the decimal divide the percentage by 100.

OpenStudy (anonymous):

Use a proportion. 75% is 75/100 well you need to find x in 24/x 75/100 = 24/x Find this by multiplying 24 by 100, then dividing your answer by 75. That will give you X.

OpenStudy (anonymous):

 what is the measure of each exterior angle of a regular triangle?